A car is traveling to the right with a speed of 2.0m/s ​ on an icy road when the brakes are applied. The car begins sliding with constant acceleration for 3.0m until it comes to a stop. How long does it take the car to slide to a stop?

Answer using a coordinate system where rightward is positive. Round the answer to two significant digits.
Answer #1

U = 2.0m/s , V = 0 (stop) , S = 3m , t =?

From V^2 - U^2 = 2aS

=) a = -4/6 = -0.667m/s^2

Now again by V-U = at

We have t = -U/a = 2/0.667 = 3s

Hence require time is t = 3.0s